True or false: density multiplied by volume equals mass.

True or false: moles multiplied by molecular weight equals grams.

True or false: moles divided by molecular weight equals grams.
Answer #1

1)

Unit of density is g/mL

Unit of volume is mL

So,

Density * volume

= g/mL * mL

= g

This is unit of mass

Answer: True

2)

Unit of mol is mol

Unit of molecular weight is g/mol

So,

Mol * molecular weight

= mol * g/mol

= g

This is unit of mass

Answer: True

3)

Unit of mol is mol

Unit of molecular weight is g/mol

So,

Mol * molecular weight

= mol / g/mol

= mol^2/g

This is not unit of mass

Answer: False